# Conditional score-based diffusion models for solving inverse problems in mechanics

### v2

@@ 4.2 Applications with experimental data
In this section, we apply the cSDMs to three inverse problems with experimental data. We train the score network on synthetic data but perform inference using experimental data. These examples help demonstrate the utility of the proposed approach to problems of practical relevance.
@@ 4.2.1 Quasi-static elastography of a circular inclusion within homogeneous media
For an elastic medium, the relation between the displacements and shear modulus is given by equilibrium and constitutive equations Eqs. 22 and 21, respectively. We also make plane stress and incompressibility assumptions. In this example, the specimen is 34.608$\times$26.297 mm2, with traction free boundaries along the left and right edges. Additionally, the top and bottom surfaces are traction free along the horizontal direction. To simulate compression of the specimen, the top and bottom edges are subject to 0.084 mm and 0.392 mm vertical displacements, respectively.
@@
Next, we use the trained score network and Annealed Langevin dynamics to sample the posterior distribution and infer the spatial distribution of the shear modulus for two test cases that were not part of the training dataset. Fig. 8 shows the results for these tests. In Fig. 8, the first and second columns show the true spatial distribution for the two test cases and the corresponding full-field measurements, respectively. We use Annealed Langevin dynamics sampling to generate 800 posterior realizations for each test case and then utilize them to estimate the pixel-wise posterior mean and standard deviation. We show the posterior statistics in the third and fourth columns of Fig. 8. Finally, the last column in Fig. 8 shows the absolute pixel-wise error between the posterior mean estimated using the conditional diffusion model and the corresponding ground truth. The RMSE between the posterior mean and the ground truth are 0.46 and 0.63 for the two test cases. We observe relatively larger uncertainty around the edges of the stiff inclusion and comparatively less uncertainty about the stiffness of the inclusion.
Figure 8: Posterior statistics estimated using the conditional diffusion model on test samples for the quasi-static elastography application
@@
Figure 9: Posterior statistics estimated using the conditional diffusion model on experimental data for the quasi-static elastography application
@@ 4.2.2 Time-dependent elastography of a dual-layer tissue-mimicking phantom
Fig. 11 shows the results for two test samples where we use synthetic measurements to infer the spatial distribution of the Young’s moduli. The first column shows the ground truth for the Young’s modulus distribution of the phantom. The second to fourth columns show three posterior realizations. The fifth and sixth columns show the pixel-wise posterior mean and standard deviation estimated using the conditional score-based diffusion model with 500 posterior realizations. The posterior mean provides a good estimate for the ground truth both in terms of the Young’s modulus for each layer and the location of the interface. Table 10 quantitatively compares the estimated posterior mean values of the Young’s modulus and the location of the interface to their corresponding true values.
@@
Figure 11: Posterior statistics estimated using the conditional diffusion model on test data for the time-dependent elastography application
@@
Figure 13: Posterior realizations (first three columns) and statistics estimated using the conditional diffusion model on experimental data for the ultrasound elastography application
Fig. 13 shows the results of the inference for the experimental data collected by Lu et al. [57] and shown in Fig. 12. This figure shows three typical samples from the posterior distribution and the pixel-wise mean and standard deviation computed using 500 realizations from the posterior distribution. The posterior mean estimate for the Young’s modulus of the top and bottom layers is 6.98 kPa and 19.73 kPa, respectively. Uniaxial quasi-static mechanical tests performed on homogenous phantoms with the same gelatin concentrations as the two layers yielded values of 2.1 $\pm$ 0.2 kPa and 20.3 $\pm$ 0.4 kPa. Thus, the cSDM estimates the modulus of the lower layer with good accuracy. However, there is a significant error in the estimate for the modulus of the top layer, which may be because this layer is relatively thin and, therefore, its contribution to the measured displacement is unlikely to be significant. The phantom was fabricated such that the interface would be at a height of 3.0 mm. The cSDM estimate for the mean interface height is 3.21 mm with an average standard deviation of 0.10 mm; the true location of the interface is close to two standard deviations of the estimated posterior mean.
@@ 4.2.3 Optical coherence elastography of tumor spheroids
Optical coherence tomography (OCT) is a technique that leverages optics, analogous to acoustics in ultrasound imaging, to produce high-resolution tomograms of tissue. A functional extension of OCT that enables imaging based on mechanical contrast is optical coherence elastography (OCE) [72], which involves capturing the phase data alongside intensity measurements of the interference of light backscattered from the specimen with light reflected from a reference mirror at varied axial depths. One variant of OCE is phase-sensitive compression OCE [73], wherein the phase shifts observed between B-scans (cross-sectional images) of the specimen as it undergoes compressive loading quantify the specimen’s deformation along the optical axis. The mechanical properties of the specimen can be inferred from these phase shifts because they encode within them information about the displacement component along the optical axis. In this study, we consider a high resolution variant of phase-sensitive compression OCE, termed mechano-microscopy [74, 75], which utilizes a high-resolution variant of OCT known as optical coherence microscopy (OCM). A conditional score-based diffusion model is used to solve the inverse problem of determining elasticity profiles, specifically the spatial distribution of the Young’s modulus ${\mathbf{X}}$ of tumor spheroids from noisy phase difference measurements ${\mathbf{Y}}$. The data for this experiment were obtained from measurements and datasets originally compiled by Foo et al. [58].
We provide a brief summary of the experimental setup here (Fig. 14) and the reader is referred to [58] for more details. In the experiment, a custom-built spectral-domain OCM system in dual-arm configuration images tumor spheroids embedded in hydrogel specimens from the top. Following the objective lens, the sample arm contains, from top to bottom: a rigid glass coverslip acting as an imaging window, a 300 $\mu$m thick sample containing tumor spheroids, a 1 mm thick compliant silicone layer, and a piezoelectric actuator to apply compressive displacement. The purpose of the compliant layer is to provide surface stress measurements for mechano-microscopy [74]; we do not utilize this data in this study. Two volumetric images at different compression levels were acquired at each location to produce a volumetric phase difference image. The plane corresponding to the central cross-section of the tumor spheroid was identified in the volumetric image and was cropped to a 256$\times$256 $\mu$m2 window around the spheroid. This two-dimensional phase difference image was used as measurement in solving the inverse problem.
@@
The specimens were prepared by encapsulating MCF7 (non-metastatic breast cancer) cells within a hydrogel matrix composed of lyophilized gelatin methacryloyl (GelMA). The cells were cultured until they formed spheroids. Each tumor spheroid contains multiple cells, and each cell domain can be broadly classified into two components: the cytoplasm and the nucleus (see Fig. 15). We take inspiration from this morphology to develop the parametric prior for this study. We refer interested readers to [76] for typical three-dimensional immuno-fluorescence images of tumor spheroids taken at different time points.
The synthetic samples, which we generate from the parametric prior distribution, represent a spheroid within a hydrogel on a 256$\times$256 $\mu$m2 domain (first column in Fig. 16). We model the spheroid as an ellipse, and straight lines subdivide the spheroid into cells. Each cell contains a circle representing a nucleus, surrounded by the cytoplasm of the cell. We treat the following parameters as truncated normal random variables: spheroid width, spheroid aspect ratio, spheroid position, number of nuclei per unit area, nuclei radii, nuclei elasticity, cytoplasm elasticity, and hydrogel elasticity. Table 11 provides details on these parameter’s distributions, which we derive from preliminary analyses of spheroid OCM images. We determine the nuclei locations (defined by their center coordinates) using a custom algorithm that iteratively accumulates nuclei coordinates by creating a list of valid coordinates for the current nucleus to be positioned, randomly selecting a coordinate from this list for the current nucleus, removing the coordinates occupied by the current nucleus from the list of valid coordinates for the next nucleus, and repeating until the required number of nuclei for the spheroid is reached or there are no more valid coordinates for the next nucleus. We constrain the sizes and positions of the nuclei to ensure a minimum spacing of 2 $\mu$m between nuclei and between nuclei and the ellipse boundary. We subsequently determine the cell boundaries using a power diagram, also known as a Laguerre–Voronoi diagram [77].
Table 11: Random variables comprising the parametric prior for the tumor spheroid application
Parameter Variable Minimum Maximum
Spheroid width ($\mu$m) $\mathcal{N}(105,8.33)$ 80 130
Spheroid height to width ratio $\mathcal{N}(0.90,0.0167)$ 0.85 0.95
Radius of nuclei ($\mu$m) $\mathcal{N}(7.5,0.5)$ 6 9
Young’s modulus of nuclei (kPa) $\mathcal{N}(32.5,5.833)$ 15 50
Young’s modulus of cytoplasm (kPa) $\mathcal{N}(10,1.67)$ 5 15
Young’s modulus of hydrogel (kPa) $\mathcal{N}(7.5,0.833)$ 5 10
Displacement applied to top edge ($\mu$m) $\mathcal{N}(-2.2,0.6)$ $-0.4$ $-4$
Fig. 15 shows the steps for generating realizations of ${\mathbf{Y}}$ from realizations of ${\mathbf{X}}$. Given a realization ${\mathbf{x}}$ of the Young’s modulus field, we simulate uniaxial compression on the specimen via finite element analysis (FEA) in ABAQUS [71] to obtain the vertical displacement field (third column in Fig. 15). To mitigate edge effects, we initially expand each sample to an 800$\times$800 $\mu\text{m}^{2}$ domain by padding with hydrogel on all four sides. We displace the top edge downward while the bottom-left corner remains fixed, and the bottom edge is constrained vertically. We sample the displacement on the top edge from a truncated normal distribution; see Table 11. The magnitude of the displacement of the top edge ensures that the corresponding strain values are within the order of a few millistrains, consistent with observations from mechano-microscopy experiments. For this analysis, given the low levels of strain involved, we assume that the specimen is linear and elastic (Eq. 21), under plane strain conditions (Eq. 22), and nearly incompressible with Poisson’s ratio uniformly set to 0.49. We develop a CAD model for the specimen (second column in Fig. 15), which we subsequently use to develop the FEA mesh geometry consisting of quadrilateral elements for the main region and triangular elements for the padded regions. Once the FEA is complete, we obtain 2D arrays of the Young’s modulus $E$ and vertical displacement $u_{2}$ by interpolating the main 256$\times$256 $\mu\text{m}^{2}$ domain window to a 256$\times$256 Cartesian grid. We convert the displacement fields to phase difference fields $\Delta\phi$ using the following equation:
$\Delta\phi=\frac{4\pi n_{r}u_{2}}{\lambda_{0}}\;,$ (25)
@@
In this manner, we initially synthesize a synthetic dataset comprising 5000 pairs of $\bm{X}$ and $\bm{Y}$ realizations. We augment this dataset sixfold by shifting the 256$\times$256 $\mu\text{m}^{2}$ image window 20 $\mu$m in each horizontal and vertical direction (left, right, up, and down), as well as applying a horizontal flip to each image. We divide the resulting dataset into a training set of 24000 images and a test set of 6000 images. As a preprocessing step, we rescale the elasticity images to enhance contrast and facilitate better training convergence. The rescaling involves the following transformation
@@
Fig. 17 shows realizations from the posterior, and its statistics, estimated using the conditional score-based diffusion model on both synthetic and experimental data. The first two columns correspond to samples from the test dataset previously unseen by the model, while columns three and four depict predictions on experimental measurements. The first row shows the measurements used for sampling using the trained score network. Rows two to four show three realizations sampled by the model for the corresponding measurements. Rows five and six show the pixel-wise posterior mean and standard deviation, respectively, estimated using 200 realizations. For the synthetic samples, we compare the estimated posterior means for the Young’s modulus distribution with the corresponding ground truth. In the case of experimental data, comparisons were made against OCM images that were co-registered during the acquisition of the phase difference images. Note that all images are presented at their true scales. Recall that the recovery of the Young’s modulus is possible up to a multiplicative constant: the hydrogel modulus $E_{\text{hydrogel}}$ in this case. We assume $E_{\text{hydrogel}}=$ 10 kPa for the test samples. For the experimental samples, $E_{\text{hydrogel}}$ was estimated from mechano-microscopy measurements [58].
@@ 5 Conclusions, outlook, and extensions
Large-scale Bayesian inference using black-box models is challenging. In this work, we explore the use of conditional score-based diffusion models for large-scale Bayesian inference. Fundamental to the score-based diffusion models is the score network that can be trained using realizations from the joint distribution between the inferred quantity and the measurements. Thus, we only need to simulate from black-box models to create the training dataset. The trained score network can be used to sample the target posterior distribution using Annealed Langevin dynamics. The score network learns the score function of the target posterior ‘averaged’ over the joint distribution. Therefore, there is no need to re-train the score function when a new measurement is made. This makes conditional score-based diffusion models particularly suitable for high-throughput applications.
@@
We apply conditional score-based diffusion models to solve several inverse problems in mechanics. We validate the proposed approach on two synthetic examples; for these problems the quality of the inference is similar to Monte Carlo simulation. Further, we demonstrate the utility of the proposed approach on three applications where experimental measurements are available. Overall, we show that conditional score-based diffusion models can tackle high-dimensional inverse problems involving complex prior distributions, different types of forward physics simulated using black-box models, complex measurement noise models and modeling errors, and various measurement modalities.
@@
Although the results in this manuscript are promising, we identify several interesting future research directions that will help with the adoption of this technology as an inference toolbox. First, the primary computational bottleneck in applying conditional score-based diffusion models, and any conditional generative algorithm for that matter, will be the cost of curating the training dataset, which requires multiple simulations of the black-box forward model. Thus, there is a need for the development of data-efficient conditional diffusion models. Moreover, training using multi-fidelity datasets promises to be another interesting research direction. Further, we consider full-field measurements in this work. Extensions to sparse measurements using masking-based training strategies [80] is another extension that will increase the utility of the proposed method. Finally, applying conditional score-based diffusion models to physics-based inverse problems that arise in different science and engineering domains will be interesting, as it will reveal insights into the behavior of conditional generative algorithms of this type.

### v3

@@ 4.2 Applications with experimental data
In this section, we apply the cSDMs to three inverse problems with experimental data. We train the score network on synthetic data but perform inference using experimental data. These examples help demonstrate the utility of the proposed approach on problems of practical relevance.
@@ 4.2.1 Quasi-static elastography of a circular inclusion within homogeneous media
For an elastic medium, the relation between the displacements and shear modulus is given by equilibrium and constitutive equations Eqs. 22 and 21, respectively. We also make plane stress and incompressibility assumptions similar to previous studies [42, 46]. In this example, the specimen is 34.608$\times$26.297 mm2, with traction free boundaries along the left and right edges. Additionally, the top and bottom surfaces are traction free along the horizontal direction. To simulate compression of the specimen, the top and bottom edges are subject to 0.084 mm and 0.392 mm vertical displacements, respectively.
Next, we use the trained score network and Annealed Langevin dynamics to sample the posterior distribution and infer the spatial distribution of the shear modulus for two test cases that were not part of the training dataset. Fig. 9 shows the results for these tests. In Fig. 9, the first and second columns show the true spatial distribution for the two test cases and the corresponding full-field measurements, respectively. We use Annealed Langevin dynamics sampling to generate 800 posterior realizations for each test case and then utilize them to estimate the pixel-wise posterior mean and standard deviation. We show the posterior statistics in the third and fourth columns of Fig. 9. Finally, the last column in Fig. 9 shows the absolute pixel-wise error between the posterior mean estimated using the cSDM and the corresponding ground truth. The RMSE between the posterior mean and the ground truth are 0.46 and 0.63 for the two test cases. We observe relatively larger uncertainty around the edges of the stiff inclusion and comparatively less uncertainty about the stiffness of the inclusion.
Figure 9: Posterior statistics estimated using the cSDM on test samples for the quasi-static elastography application
Figure 10: Posterior statistics estimated using the cSDM on experimental data for the quasi-static elastography application
@@ 4.2.2 Time-dependent elastography of a dual-layer tissue-mimicking phantom
Figure 12: Posterior statistics estimated using the cSDM on test data for the time-dependent elastography application
Fig. 12 shows the results for two test samples where we use synthetic measurements to infer the spatial distribution of the Young’s moduli. The first column shows the ground truth for the Young’s modulus distribution of the phantom. The second to fourth columns show three posterior realizations. The fifth and sixth columns show the pixel-wise posterior mean and standard deviation estimated using the cSDM with 500 posterior realizations. The posterior mean provides a good estimate for the ground truth both in terms of the Young’s modulus for each layer and the location of the interface. Table 12 quantitatively compares the estimated posterior mean values of the Young’s modulus and the location of the interface to their corresponding true values.
@@
Figure 14: Posterior realizations (first three columns) and statistics estimated using the cSDM on experimental data for the ultrasound elastography application
Fig. 14 shows the results of the inference for the experimental data collected by Lu et al. [61] and shown in Fig. 13. This figure shows three typical samples from the posterior distribution and the pixel-wise mean and standard deviation computed using 500 realizations from the posterior distribution. The posterior mean estimate for the Young’s modulus of the top and bottom layers is 6.98 kPa and 19.73 kPa, respectively. Uniaxial quasi-static mechanical tests performed on homogenous phantoms with the same gelatin concentrations as the two layers yielded values of 2.1 $\pm$ 0.2 kPa and 20.3 $\pm$ 0.4 kPa. Thus, the cSDM estimates the modulus of the lower layer with good accuracy. However, there is a significant error in the estimate for the modulus of the top layer. We observed a similar overestimation in the Young’s modulus of a thin top layer on a test example similar to the phantom but not part of the training dataset. There are two plausible sources for this error. The overestimation of the Young’s modulus of a thin top layer may be because its contribution to the measured displacement is unlikely to be significant. Hence, the probabilistic model for modeling errors, which includes discrepancies in the forward physics and measurement noise, may require revision to account for such atypical cases. Alternatively, the erroneous prediction may be due to a lack of similar examples in the training dataset and the score network not generalizing well to such cases. The phantom was fabricated such that the interface would be at a height of 3.0 mm. The estimate for the mean interface height obtained using the cSDM is 3.21 mm with an average standard deviation of 0.10 mm; the true location of the interface is close to two standard deviations of the estimated posterior mean.
@@ 4.2.3 Optical coherence elastography of tumor spheroids
The final application concerns OCE–an optical coherence tomography (OCT) method that enables mechanical contrast-based imaging using the phase data of backscattered light from a specimen undergoing compression [76, 77]. More specifically, this application considers mechano-microscopy [78, 79] of tumor spheroids. Mechano-microscopy is a high resolution variant of phase-sensitive compression OCE, which utilizes a high-resolution variant of OCT known as optical coherence microscopy (OCM). Tumor spheroids refer to a collection of multiple tumor cells. We provide a short background on OCE in Section C1 and refer interested readers to [80] for typical three-dimensional immuno-fluorescence images of tumor spheroids taken at different time points. In this application, we use a cSDM to infer the spatial distribution of the Young’s modulus ${\mathbf{X}}$ of tumor spheroids from noisy phase difference measurements ${\mathbf{Y}}$, which encode information regarding the specimen’s displacement along the optical axis. Fig. 15 provides a schematic of the experimental setup. We relegate additional details regarding the physical experiment (Section C2) and specimen preparation process (Section C3) to Appendix C. The data for this experiment were obtained from measurements and datasets originally compiled by Foo et al. [62].
@@
The synthetic samples, which we generate from the parametric prior distribution, represent a spheroid within a hydrogel on a 256$\times$256 $\mu$m2 domain (first column in Fig. 16). We describe the parametric prior for this application in Section C4. The parametric prior is highly complex, involving a custom nuclei placement algorithm and Voronoi tessellations [81]. We discretize the Young’s modulus field on a 256$\times$256 Cartesian grid over the specimen. We record noisy measurements of the wrapped phase difference field over the same grid. Fig. 16 shows the steps for generating realizations of ${\mathbf{Y}}$ from realizations of ${\mathbf{X}}$. We provide details of the forward model in Section C5, which involves developing a CAD model for the specimen, finite element analysis (FEA) using ABAQUS [75], and a non-additive non-Gaussian measurement noise model (Eq. 41). To be precise, every realization of ${\mathbf{X}}$ is a Young’s modulus field, and the corresponding realization of $Y$ is the wrapped and noisy phase difference field. Also, ${n_{\mathcal{X}}}={n_{\mathcal{Y}}}=$ 256$\times$256 for this application.
We initially synthesize a synthetic dataset comprising 5000 pairs of $\bm{X}$ and $\bm{Y}$ realizations. We augment this dataset sixfold by shifting the 256$\times$256 $\mu\text{m}^{2}$ image window 20 $\mu$m in each horizontal and vertical direction (left, right, up, and down), as well as applying a horizontal flip to each image. We divide the resulting dataset into a training set of 24000 images and a test set of 6000 images. As a preprocessing step, we rescale the elasticity images to enhance contrast and facilitate better training convergence. The rescaling involves the following transformation
@@
Fig. 18 shows realizations from the posterior, and its statistics, estimated using the cSDM on both synthetic and experimental data. The first two columns correspond to samples from the test dataset previously unseen by the model, while columns three and four depict predictions on experimental measurements. The first row shows the measurements used for sampling using the trained score network. Rows two to four show three realizations sampled by the model for the corresponding measurements. Rows five and six show the pixel-wise posterior mean and standard deviation, respectively, estimated using 200 realizations. For the synthetic samples, we compare the estimated posterior means for the Young’s modulus distribution with the corresponding ground truth. In the case of experimental data, comparisons were made against OCM images that were co-registered during the acquisition of the phase difference images. Note that all images are presented at their true scales. Recall that the recovery of the Young’s modulus is possible up to a multiplicative constant: the hydrogel modulus $E_{\text{hydrogel}}$ in this case. We assume $E_{\text{hydrogel}}=$ 10 kPa for the test samples. For the experimental samples, $E_{\text{hydrogel}}$ was estimated from mechano-microscopy measurements [62].
@@ 5 Conclusions, outlook, and extensions
Large-scale Bayesian inference using black-box models is challenging. In this work, we explore the use of cSDMs for large-scale Bayesian inference. Fundamental to the score-based diffusion models is the score network that can be trained using realizations from the joint distribution between the inferred quantity and the measurements. Thus, we only need to simulate from black-box models to create the training dataset. The trained score network can be used to sample the target posterior distribution using Annealed Langevin dynamics. The score network learns the score function of the target posterior ‘averaged’ over the joint distribution. Therefore, there is no need to re-train the score function when a new measurement is made. This makes cSDMs particularly suitable for high-throughput applications.
@@
We apply cSDMs to solve several inverse problems in mechanics. We validate the proposed approach on two synthetic examples; for these problems the quality of the inference is similar to MCS. Further, we demonstrate the utility of the proposed approach on three applications where experimental measurements are available. Overall, we show that cSDMs can tackle high-dimensional inverse problems involving complex prior distributions, different types of forward physics simulated using black-box models, complex measurement noise models and modeling errors, and various measurement modalities.
@@
Although the results in this manuscript are promising, we identify several interesting future research directions that will help with the adoption of this technology as an inference toolbox. First, the primary computational bottleneck in applying cSDMs, and any conditional generative algorithm for that matter, will be the cost of curating the training dataset, which requires multiple simulations of the black-box forward model. Thus, there is a need for the development of data-efficient conditional diffusion models. Moreover, training using multi-fidelity datasets promises to be another interesting research direction. Further, we consider full-field measurements in this work. Extensions to sparse measurements using masking-based training strategies [84] is another extension that will increase the utility of the proposed method. We also assume that models for the measurement noise and modeling errors are known because we need to sample from them when generating realizations of the observation ${\mathbf{Y}}$ from corresponding realizations of ${\mathbf{X}}$. However, complete knowledge of the measurement noise and/or modeling errors may not available a priori. In such cases, the inverse problem may require hyperpriors over parameters of the noise or error models, or require careful considerations of epistemic uncertainty surrounding them. Extending the capabilities of the current framework to handling hyperpriors or epistemic uncertainty is yet another interesting future research direction.
Finally, applying cSDMs to physics-based inverse problems that arise in different science and engineering domains will be interesting, as it will reveal insights into the behavior of conditional generative algorithms of this type. Quantifying the reliability of the inference results obtained using conditional generative models is also important. Evaluating the trained conditional generative model using posterior predictive checks and cross validation is necessary [85]. However, the required auxiliary datasets to perform posterior predictive checks or the computational budget necessary for cross validation may be unavailable. Alternative approaches to assessing estimates of aleatoric uncertainties–primary sources of uncertainty in inverse problems of the type we consider–similar to metrics to evaluate epistemic uncertainty that stems from neural network models [86] are also needed.